# Optimal approximation of infinite-dimensional holomorphic functions II: recovery from i.i.d. pointwise samples

### v2

@@ 5.4 Bounds for Jacobi polynomial coefficients of holomorphic functions and an abstract summability lemma
$\leq\frac{(\tilde{c}(\nu_{i}+1)+1)^{\tilde{\gamma}}}{(\tilde{c}\nu_{i}+1)^{\tilde{\gamma}}}\left(\beta+\frac{\varepsilon(\nu_{i}+1)}{4\tilde{b}_{i}({\|\bm{\nu}_{F}\|}_{1}+1)}\right)^{-1}\left(\dfrac{\|\bm{\nu}_{F}\|_{1}}{\|\bm{\nu}_{F}\|_{1}+1}\right)^{-\|\bm{\nu}_{F}\|_{1}}.$
Here, we used the fact that $\frac{a+1}{b+1}\geq\frac{a}{b}$ for all $b\geq a>0$. Now observe that
$\dfrac{(\tilde{c}(\nu_{i}+1)+1)}{{(\tilde{c}\nu_{i}+1)}}=\dfrac{(\nu_{i}+1)(\tilde{c}+\frac{1}{(\nu_{i}+1)})}{\nu_{i}(\tilde{c}+\frac{1}{\nu_{i}})}\leq 2.$
Therefore, we get
@@
We now show that the sequence $\widetilde{B}(\bm{\nu})$ is anchored, i.e., $\widetilde{B}(\bm{e}_{j})\leq\widetilde{B}(\bm{e}_{i})$ for all $i,j\in\mathbb{N}$ with $i\leq j$. We proceed as [5, Thm. 3.33, Step 6]. Observe that
@@
Since $D\geq 1$ by definition, this is guaranteed by choosing $\beta\geq(\tilde{c}+1)^{\tilde{\gamma}}/\eta$. Combining this with (5.30) and the definition of $\eta$, we now suppose that
$\beta\geq(\max\{\tilde{c},1\}+1)^{\tilde{\gamma}}\max\{\mathrm{e},2\tilde{\kappa}/(\tilde{\kappa}+1)\}.$ (5.31)
With any such $\beta$, we conclude that $\widetilde{B}(\bm{\nu})$ is monotonically decreasing and anchored.
@@
$\sum_{j>d}\tilde{b}_{j}<\frac{\varepsilon}{4\mathrm{e}}.$
This completes the proof of (5.17).
@@
Therefore, (5.16) follows immediately from (5.18). Hence, it remains to show (5.18)
@@
where $\eta=\frac{1+\tilde{\kappa}}{2\tilde{\kappa}}\geq 1$, $D=D(\tilde{c},\tilde{\gamma},\tilde{\kappa})$ is any constant such that (5.29) holds, $\tilde{\bm{g}}(\bm{b},\varepsilon)$ is as in (5.35) and $d\in\mathbb{N}$ is sufficiently large so that (5.27) holds for any $\beta$ satisfying (5.31).
@@
$C_{2}(\bm{b},r,\varepsilon,c,\gamma,\xi)\leq(C_{3}(r,\varepsilon,c,\gamma,\xi))^{d(p,\varepsilon,c,\gamma,\xi)}\cdot C_{4}(p,r,\varepsilon,c,\gamma,\xi)=\widetilde{C}(p,r,\varepsilon,c,\gamma,\xi)$
Since $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$, ${\|\bm{b}\|}_{p,\mathsf{M}}\leq 1$ was arbitrary, we deduce that (5.18) holds, and therefore (5.16) as well. This completes the proof. ∎
5.5 Proofs of Theorems 5.1 and 5.2
We now prove Theorems 5.1 and 5.2. As noted, this involves constructing certain surrogate sequences, then asserting their summability, and finally, applying appropriate variants of Stechkin’s inequality.
@@
where $C(\bm{b},p,\overline{c},\overline{\gamma})$ is the constant from (5.15) with $c$ and $\gamma$ replaced by $\bar{c}$ and $\bar{\gamma}$, respectively, and the above choices of $\varepsilon$ and $\xi$. Having shown weighted summability, we are now ready to prove (5.9) and the assertion below it that the set $S$ can be chosen in a nested manner. We do this using the weighted Stechkin’s inequality (see [5, Lem. 3.12]) and, for the latter, by inspecting its proof. The proof of this inequality constructs the set $S$ as follows. Let $\pi:\mathcal{F}\rightarrow\mathbb{N}$ be a bijection that gives a nonincreasing rearrangement of the sequence $(d_{\bm{\nu}}/u_{\bm{\nu}})_{\bm{\nu}\in\mathcal{F}}$ in absolute value, i.e., $|d_{\pi(1)}|/u_{\pi(1)}\geq|d_{\pi(2)}|/u_{\pi(2)}\geq\cdots$. Then let $M\in\mathbb{N}\cup\{\infty\}$ be the maximal integer such that $\sum^{M}_{j=1}u^{2}_{\pi(j)}\leq k$ and define $S=\{\pi(1),\ldots,\pi(M)\}$. The remainder of the proof of [5, Lem. 3.12] goes on to show that this choice of $S$ yields the bound
@@
Observe that this set $S$ depends on the intrinsic weights $\bm{u}$ and the sequence $\bm{d}$ only. The former depend on $\bm{\alpha}$ and $\bm{\beta}$, while the latter depends on $\bm{b}$. Hence $S=S(\bm{b},\bm{\alpha},\bm{\beta})$, as required. Moreover, it is clear from the construction that $S$ can be chosen so that the nestedness property holds. Using this, we deduce that
@@
${\|\bar{\bm{d}}\|}_{p,\mathsf{A}}\leq C(\bm{b},p,\overline{c},\overline{\gamma}).$
@@
The bound (5.10) now follows, since $\bar{c}=\bar{c}(q,\tau)$ and $\bar{\gamma}=\bar{\gamma}(q,\tau)$. Observe also that the set $S$ is anchored by construction and depends on $\bar{\bm{d}}$. Since $\bar{\bm{d}}$ depends on $\bm{b}$, $\bm{\alpha}$, $\bm{\beta}$ and $q$ only, we deduce that $S$ has the desired form. Finally, the nestedness property follows immediately, due to the construction of $S$ as an anchored set containing $s$ largest entries of $\tilde{\bm{d}}$. This completes the proof of part (a).
@@ 6.1 Reconstruction map: known anisotropy
We commence with the former. Our first task is to define the reconstruction map $\mathcal{T}$. Let $\mathcal{U}=[-1,1]^{\mathbb{N}}$, $\mathcal{V}$ be a Hilbert space, $\bm{\alpha},\bm{\beta}$ satisfy Assumption 4.1 and $\varrho=\varrho_{\bm{\alpha},\bm{\beta}}$ be as in (4.3). Let $m\geq 3$, $10<p<$ 1, $10<\epsilon<$ 1 and $\bm{b}=(b_{i})_{i\in\mathbb{N}}\in\ell^{p}(\mathbb{N})$ with $\bm{b}\geq{\bm{0}}$ be given, and define
@@
Let $\bm{\alpha}_{\pi}=\pi(\bm{\alpha})$ and $\bm{\beta}_{\pi}=\pi(\bm{\beta})$ and consider the tensor-product Jacobi polynomials (4.5) corresponding to $\bm{\alpha}_{\pi}$ and $\bm{\beta}_{\pi}$ and let the weights $\bm{u}_{\pi}$ be as in (5.8), also for $\bm{\alpha}_{\pi}$ and $\bm{\beta}_{\pi}$. Theorem 5.1, part (a) with $q=2$ now implies the existence of an index set $S_{\pi}=S_{\pi}(\bm{b}_{\pi},\bm{\alpha}_{\pi},\bm{\beta}_{\pi})$ with $|S_{\pi}|\leq|S_{\pi}|_{\bm{u}_{\pi}}\leq k$ such that
@@
where $g_{S_{\pi}}$ is as in (5.3) with the Jacobi polynomials corresponding to $\bm{\alpha}_{\pi}$ and $\bm{\beta}_{\pi}$ and $\varrho_{\pi}=\varrho_{\bm{\alpha}_{\pi},\bm{\beta}_{\pi}}$ is as in (4.3). Here, we also used Parseval’s identity and the fact that ${\left\|\cdot\right\|}_{2,\bm{u};\mathcal{V}}={\left\|\cdot\right\|}_{2;\mathcal{V}}$. Following this, we now define the set $S=S(\bm{b},\bm{\alpha},\bm{\beta})$ by
@@
Note that reason for introducing the permutation $\pi$ is to obtain a bound (6.2) depending on the monotonic sequence $\bm{b}_{\pi}$. If we worked directly with $\bm{\alpha}$ and $\bm{\beta}$, it would only be possible to derive a bound for for $\overline{\theta_{m}}(p,\mathsf{M})$ in Theorem 4.3, as opposed to one for $\overline{\theta_{m}}(p)$.
Having defined $S$, we now construct the reconstruction map. Analogously to [3, Eqn. (5.17)], given any sample points $\{\bm{y}_{i}\}^{m}_{i=1}\subset\mathcal{U}$ we define the matrix $\bm{A}$ by
@@
The map $\mathcal{T}$ is similar to that used in [3, Thm. 4.4]. However, there are some key differences. First, we now consider arbitrary Jacobi polynomials, which may also have different parameters $\alpha_{i}$, $\beta_{i}$ in each variable. The latter necessitates the use of the permutation $\pi$ in the construction of $\mathcal{T}$ and a more involved argument in the proof below. Conversely, [3, Thm. 4.4] only considers tensor-product Chebyshev or Legendre polynomials. Second, in Theorem 4.5 we establish a stronger uniform guarantee that holds simultaneously for all functions $f\in\mathcal{H}(\bm{b})$, whereas [3, Thm. 4.4] is nonuniform and therefore weaker. Its proof consequently requires a different approach to that of [3, Thm. 4.4]. This is based on ideas found in [25, 26], and involves expanding the error in a sequence of nested index sets. See Remark 6.2 for some further discussion on this point.
@@
$\bm{z}^{*}\bm{X}_{i}\bm{z}=\frac{1}{m}\left|\sum_{\bm{\nu}\in S}z_{\bm{\nu}}\Psi_{\bm{\nu}}(\bm{y}_{i})\right|^{2}\leq\frac{|S|_{\bm{u}}}{m}{\|\bm{z}\|}^{2}_{2}.$
Therefore, $\lambda_{\max}(\bm{X}_{i})\leq|S|_{\bm{u}}/m\leq k/m$ almost surely. Hence, an application of the matrix Chernoff bound gives that
@@
Now consider the matrices $\bm{A}^{(j)}$, $j=1,\ldots,r$. By the same arguments, except with the bound $k/m$ replaced by $k_{j}/m$, we deduce that
@@
    The proofs of [3, Thm. 4.4] and Theorem 4.5 diverge in how they bound the discrete error (6.9). The former uses the fact that ${\|f-f_{S}\|}^{2}_{\mathsf{disc}}$ is a sum of independent random variables with mean ${\|f-f_{S}\|}^{2}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}$ to provide a nonuniform bound via Bernstein’s inequality. Conversely, the proof of Theorem 4.5 uses the splitting (6.10) and repeated applications of the matrix Chernoff inequality to derive a stronger uniform bound.

### v3

@@ 5.5 Bounds for Jacobi polynomial coefficients of holomorphic functions and an abstract summability lemma
$\dfrac{\widetilde{B}_{F}(\bm{\nu}+\bm{e}_{i})}{\widetilde{B}_{F}(\bm{\nu})}$ $=\frac{(\tilde{c}(\nu_{i}+1)+1)^{\tilde{\gamma}}}{(\tilde{c}\nu_{i}+1)^{\tilde{\gamma}}}\frac{\left(\beta+\frac{\varepsilon(\nu_{i}+1)}{4\tilde{b}_{i}({\|\bm{\nu}_{F}\|}_{1}+1)}\right)^{-(\nu_{i}+1)}}{\left(\beta+\frac{\varepsilon\nu_{i}}{4\tilde{b}_{i}{\|\bm{\nu}_{F}\|}_{1}}\right)^{-\nu_{i}}}\prod_{j\in\mathrm{supp}(\bm{\nu}_{F})\backslash\{i\}}\left(\frac{\beta+\frac{\varepsilon\nu_{j}}{4\tilde{b}_{j}({\|\bm{\nu}_{F}\|}_{1}+1)}}{\beta+\frac{\varepsilon\nu_{j}}{4\tilde{b}_{j}{\|\bm{\nu}_{F}\|}_{1}}}\right)^{-\nu_{j}}.$
Consider the three terms on the right hand side separately. For the first, we observe that
$\dfrac{(\tilde{c}(\nu_{i}+1)+1)^{\tilde{\gamma}}}{(\tilde{c}\nu_{i}+1)^{\tilde{\gamma}}}=\left(\dfrac{(\nu_{i}+1)(\tilde{c}+\frac{1}{(\nu_{i}+1)})}{\nu_{i}(\tilde{c}+\frac{1}{\nu_{i}})}\right)^{\tilde{\gamma}}\leq 2^{\tilde{\gamma}}.$
For the second, we use the inequality $\frac{a+c}{b+c}\geq\frac{a}{b}$ for all $b\geq a>0$ and $c\geq 0$. Since $i>d$, and therefore $\nu_{i}\leq{\|\bm{\nu}_{F}\|}_{1}$, an application of this inequality with $c=1$ gives
$\frac{\left(\beta+\frac{\varepsilon(\nu_{i}+1)}{4\tilde{b}_{i}({\|\bm{\nu}_{F}\|}_{1}+1)}\right)^{-(\nu_{i}+1)}}{\left(\beta+\frac{\varepsilon\nu_{i}}{4\tilde{b}_{i}{\|\bm{\nu}_{F}\|}_{1}}\right)^{-\nu_{i}}}\leq\left(\beta+\frac{\varepsilon(\nu_{i}+1)}{4\tilde{b}_{i}({\|\bm{\nu}_{F}\|}_{1}+1)}\right)^{-1}\frac{\left(\beta+\frac{\varepsilon\nu_{i}}{4\tilde{b}_{i}{\|\bm{\nu}_{F}\|}_{1}}\right)^{-\nu_{i}}}{\left(\beta+\frac{\varepsilon\nu_{i}}{4\tilde{b}_{i}{\|\bm{\nu}_{F}\|}_{1}}\right)^{-\nu_{i}}}\leq\beta^{-1}.$
For the third term, we use the same inequality again, except with $c=\beta$. This gives
$\prod_{j\in\mathrm{supp}(\bm{\nu}_{F})\backslash\{i\}}\left(\frac{\beta+\frac{\varepsilon\nu_{j}}{4\tilde{b}_{j}({\|\bm{\nu}_{F}\|}_{1}+1)}}{\beta+\frac{\varepsilon\nu_{j}}{4\tilde{b}_{j}{\|\bm{\nu}_{F}\|}_{1}}}\right)^{-\nu_{j}}\leq\prod_{j\in\mathrm{supp}(\bm{\nu}_{F})\backslash\{i\}}\left(\dfrac{\|\bm{\nu}_{F}\|_{1}}{\|\bm{\nu}_{F}\|_{1}+1}\right)^{-\nu_{j}}\leq\left(\dfrac{\|\bm{\nu}_{F}\|_{1}}{\|\bm{\nu}_{F}\|_{1}+1}\right)^{-\|\bm{\nu}_{F}\|_{1}}\leq\mathrm{e}.$
Here, in the last step, we used the elementary inequality $(1+1/n)^{n}\leq\mathrm{e}$, $\forall n\in\mathbb{N}$. Combining this with the other two bounds, we deduce that
@@
This shows that $\widetilde{B}_{F}(\bm{\nu})$ is monotonically nonincreasing, and therefore so is $\widetilde{B}(\bm{\nu})$.
We now show that this sequence is anchored, i.e., $\widetilde{B}(\bm{e}_{j})\leq\widetilde{B}(\bm{e}_{i})$ for all $i,j\in\mathbb{N}$ with $i\leq j$. We proceed as [5, Thm. 3.33, Step 6]. Observe that
@@
Since $D\geq 1$ by definition, this is guaranteed by choosing $\beta\geq(\tilde{c}+1)^{\tilde{\gamma}}/\eta$. However, $1/\eta=2\tilde{\kappa}/(1+\tilde{\kappa})\leq 2$. So this is implied by (5.32). With any such $\beta$, we conclude that $\widetilde{B}(\bm{\nu})$ is both monotonically decreasing and anchored.
@@
$\sum_{j>d}\tilde{b}_{j}<\frac{\varepsilon}{4\mathrm{e}}$
holds, in addition to (5.29). This completes the proof of (5.19).
@@
This implies that (5.18) is a consequence of (5.20), and therefore it remains to show (5.20).
@@
where $\eta=\frac{1+\tilde{\kappa}}{2\tilde{\kappa}}\geq 1$, $D=D(\tilde{c},\tilde{\gamma},\tilde{\kappa})$ is any constant such that (5.31) holds, $\tilde{\bm{g}}(\bm{b},\varepsilon)$ is as in (5.36) and $d\in\mathbb{N}$ is sufficiently large so that (5.29) holds for $\beta$ satisfying (5.32).
@@
$C_{2}(\bm{b},r,\varepsilon,c,\gamma,\xi)\leq(C_{3}(r,\varepsilon,c,\gamma,\xi))^{d(p,\varepsilon,c,\gamma,\xi)}\cdot C_{4}(p,r,\varepsilon,c,\gamma,\xi)=\widetilde{C}(p,r,\varepsilon,c,\gamma,\xi).$
Since $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$, ${\|\bm{b}\|}_{p,\mathsf{M}}\leq 1$, was arbitrary, we deduce that (5.20) holds, and therefore (5.18) as well. This completes the proof. ∎
5.6 Proofs of Theorems 5.2 and 5.3
We now prove Theorems 5.2 and 5.3. As noted, this involves constructing certain surrogate sequences, then asserting their summability, and finally, applying appropriate variants of Stechkin’s inequality.
@@
where $C(\bm{b},p,\overline{c},\overline{\gamma})$ is the constant from (5.17) with $c$ and $\gamma$ replaced by $\bar{c}$ and $\bar{\gamma}$, respectively, and the above choices of $\varepsilon$ and $\xi$. Having shown weighted summability, we are now ready to prove (5.11) and the assertion below it that the set $S$ can be chosen in a nested manner. We do this using the weighted Stechkin’s inequality (see [5, Lem. 3.12]) and, for the latter, by inspecting its proof. The proof of this inequality constructs the set $S$ as follows. Let $\pi:\mathcal{F}\rightarrow\mathbb{N}$ be a bijection that gives a nonincreasing rearrangement of the sequence $(d_{\bm{\nu}}/u_{\bm{\nu}})_{\bm{\nu}\in\mathcal{F}}$ in absolute value, i.e., $|d_{\pi(1)}|/u_{\pi(1)}\geq|d_{\pi(2)}|/u_{\pi(2)}\geq\cdots$. Let $k>0$ be as in the theorem statement, $M\in\mathbb{N}\cup\{\infty\}$ be the maximal integer such that $\sum^{M}_{j=1}u^{2}_{\pi(j)}\leq k$ and define $S=\{\pi(1),\ldots,\pi(M)\}$. The remainder of the proof of [5, Lem. 3.12] goes on to show that this choice of $S$ yields the bound
@@
Here $\sigma_{k}({\bm{d}})_{q,\bm{u}}$ denotes the weighted best $(k,\bm{u})$-term approximation error in the case $\mathcal{V}=\mathbb{R}$. See (5.7). Observe that the set $S$ depends on $k$, the intrinsic weights $\bm{u}$ and the sequence $\bm{d}$ only. The former depend on $\bm{\alpha}$ and $\bm{\beta}$, while the latter depends on $\bm{b}$. Hence $S=S(k,\bm{b},\bm{\alpha},\bm{\beta})$, as required. Moreover, it is clear from the construction that $S$ can be chosen so that the nestedness property holds. Using this, we deduce that
@@
${\|\bar{\bm{d}}\|}_{p,\mathsf{A}}\leq C(\bm{b},p,\overline{c},\overline{\gamma}),$
@@
The bound (5.12) now follows, since $\bar{c}=\bar{c}(q,\tau)$ and $\bar{\gamma}=\bar{\gamma}(q,\tau)$. Observe also that the set $S$ is anchored by construction and depends on $s$ and $\bar{\bm{d}}$. Since $\bar{\bm{d}}$ depends on $\bm{b}$, $\bm{\alpha}$, $\bm{\beta}$ and $q$ only, we deduce that $S$ has the desired form. Finally, the nestedness property follows immediately, due to the construction of $S$ as an anchored set containing $s$ largest entries of $\tilde{\bm{d}}$. This completes the proof of part (a).
@@ 6.1 Reconstruction map: known anisotropy
We commence with Theorem 4.5. Our first task is to define the reconstruction map $\mathcal{T}$. Let $\mathcal{U}=[-1,1]^{\mathbb{N}}$, $\mathcal{V}$ be a Hilbert space, $\bm{\alpha},\bm{\beta}$ satisfy Assumption 4.1 and $\varrho=\varrho_{\bm{\alpha},\bm{\beta}}$ be as in (4.3). Let $m\geq 3$, $10<p<$ 1, $10<\epsilon<$ 1 and $\bm{b}=(b_{i})_{i\in\mathbb{N}}\in\ell^{p}(\mathbb{N})$ with $\bm{b}\geq\bm{0}$ be given, and define
@@
Let $\bm{\alpha}_{\pi}=\pi(\bm{\alpha})$ and $\bm{\beta}_{\pi}=\pi(\bm{\beta})$ and consider the tensor-product Jacobi polynomials (5.2) corresponding to $\bm{\alpha}_{\pi}$ and $\bm{\beta}_{\pi}$ and let the weights $\bm{u}_{\pi}$ be as in (5.10), also for $\bm{\alpha}_{\pi}$ and $\bm{\beta}_{\pi}$. Theorem 5.2, part (a) with $q=2$ now implies the existence of an index set $S_{\pi}=S_{\pi}(k,\bm{b}_{\pi},\bm{\alpha}_{\pi},\bm{\beta}_{\pi})$ with $|S_{\pi}|\leq|S_{\pi}|_{\bm{u}_{\pi}}\leq k$ such that
@@
where $g_{S_{\pi}}$ is as in (5.5) with the Jacobi polynomials corresponding to $\bm{\alpha}_{\pi}$ and $\bm{\beta}_{\pi}$ and $\varrho_{\pi}=\varrho_{\bm{\alpha}_{\pi},\bm{\beta}_{\pi}}$ is as in (4.3). Here, we also used Parseval’s identity and the fact that ${\left\|\cdot\right\|}_{2,\bm{u};\mathcal{V}}={\left\|\cdot\right\|}_{2;\mathcal{V}}$. Following this, we now define the set $S=S(k,\bm{b},\bm{\alpha},\bm{\beta})$ by
@@
Note that the reason for introducing the permutation $\pi$ is to obtain a bound (6.2) depending on the monotonic sequence $\bm{b}_{\pi}$. If we worked directly with $\bm{\alpha}$ and $\bm{\beta}$, it would only be possible to derive a bound for $\overline{\theta_{m}}(p,\mathsf{M})$ in Theorem 4.2, as opposed to one for $\overline{\theta_{m}}(p)$.
Having defined $S$, we next construct the reconstruction map. Analogously to [4, Eqn. (5.17)], given any sample points $\{\bm{y}_{i}\}^{m}_{i=1}\subset\mathcal{U}$ we define the matrix $\bm{A}$ by
@@
The map $\mathcal{T}$ is similar to that used in [4, Thm. 4.4]. However, there are some key differences. First, we now consider arbitrary Jacobi polynomials, which may also have different parameters $\alpha_{i}$, $\beta_{i}$ in each variable. The latter necessitates the use of the permutation $\pi$ in the construction of $\mathcal{T}$ and a more involved argument in the proof below. Conversely, [4, Thm. 4.4] only considers tensor-product Chebyshev or Legendre polynomials. Second, in Theorem 4.5 we establish a stronger uniform guarantee that holds simultaneously for all functions $f\in\mathcal{H}(\bm{b})$, whereas [4, Thm. 4.4] is nonuniform and therefore weaker. Its proof consequently requires a different approach to that of [4, Thm. 4.4]. This is based on ideas found in [25, 26], and involves expanding the error in a sequence of nested index sets. See Remark 6.2 for some further discussion on this point.
@@
$\bm{z}^{*}\bm{X}_{i}\bm{z}=\frac{1}{m}\left|\sum_{\bm{\nu}\in S}z_{\bm{\nu}}\Psi_{\bm{\nu}}(\bm{y}_{i})\right|^{2}\leq\frac{|S|_{\bm{u}}}{m}{\|\bm{z}\|}^{2}_{2}$
almost surely. Therefore, $\lambda_{\max}(\bm{X}_{i})\leq|S|_{\bm{u}}/m\leq k/m$ almost surely. An application of the matrix Chernoff bound now gives that
@@
Next, we consider the matrices $\bm{A}^{(j)}$, $j=1,2,\ldots$. By the same arguments, except with the bound $k/m$ replaced by $k_{j}/m$, we deduce that
@@
    The proofs of [4, Thm. 4.4] and Theorem 4.5 diverge in how they bound the discrete error in (6.9). The former uses the fact that ${\|f-f_{S}\|}^{2}_{\mathsf{disc}}$ is a sum of independent random variables with mean ${\|f-f_{S}\|}^{2}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}$ to provide a nonuniform bound via Bernstein’s inequality. Conversely, the proof of Theorem 4.5 uses the splitting (6.10) and repeated applications of the matrix Chernoff inequality to derive a stronger uniform bound.